Stan is carrying a load of 50 boxes of books in his truck. Some of the boxes weigh 20 pounds each, and the rest of the boxes weigh 10 pounds each. If all the boxes weigh a total of 900 pounds, how many 20 pound boxes are in Stan's load?
Answer by NancyLam(35)

let x= boxes weigh 20 pounds each , y = boxes weigh 10 pounds each,
20x+10y=900 pounds
x+y=50 boxes
y = 50-x
20x +10(50-x)=900pounds (subtitution)
20x+500-10x =900p
10x = 900-500
x=400/10
x =40
Subtitution x=40 for find y
40+y=50
y=50-40
y=10
Check: 20(40)+10(10)=
800 + 100=900pound
20 pouns have 40 soxes
10 pouns have 10 soxes